# WPO: Enhancing RLHF with Weighted Preference Optimization

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a promising solution to align large language models (LLMs) more closely with human values. Off-policy preference optimization, where the preference data is obtained from other models, is widely adopted due to its cost efficiency and scalability. However, off-policy preference optimization often suffers from a distributional gap between the policy used for data collection and the target policy, leading to suboptimal optimization. In this paper, we propose a novel strategy to mitigate this problem by simulating on-policy learning with off-policy preference data. Our Weighted Preference Optimization (WPO) method adapts off-policy data to resemble on-policy data more closely by reweighting preference pairs according to their probability under the current policy. This method not only addresses the distributional gap problem but also enhances the optimization process without incurring additional costs. We validate our method on instruction following benchmarks including Alpaca Eval 2 and MT-bench. WPO not only outperforms Direct Preference Optimization (DPO) by up to 5.6% on Alpaca Eval 2 but also establishes a remarkable length-controlled winning rate against GPT-4-turbo of 48.6% based on Llama-3-8B-Instruct, making it the strongest 8B model on the leaderboard. We will release the code and models at https://github.com/wzhouad/WPO.
@@ 1 Introduction
We evaluate WPO on instruction following benchmarks, including Alpaca Eval 2 Li et al. (2023) and MT-bench Zheng et al. (2023). In the off-policy setting based on Ultrafeedback Cui et al. (2023), WPO improves the length-controlled winning rate against GPT-4-turbo on Alpaca Eval 2 by up to 14.9% over SFT model, outperforming DPO by up to 5.6%. Particularly, in the hybrid RL setting where the off-policy preference data is further enriched with on-policy outputs, WPO (Figure 1) achieves a new SOTA length-controlled winning rate of 48.6% on Alpaca Eval 2, making it the strongest 8B model to date. Additionally, we find that WPO can be integrated into other loss functions for preference optimization and shows consistent improvements. Furthermore, we systematically compare the model performance in different RL settings. Our analysis reveals that the hybrid setting, which utilizes both on-policy and off-policy preference data, achieves the best results, and on-policy, dispreferred data is more important for preference optimization.
@@ 3.3 Weight Alignment
The objective of our weighting strategy is to simulate on-policy RL, where different outputs are weighed by their alignment with on-policy behavior. Therefore, after the outputs are sampled using the current policy model, their weights $w(x,y)$ should be uniform, and the weights of outputs from other models should be smaller. However, our current strategy assigns different weights to outputs sampled using the policy model, due to the fact that LLMs exhibit different levels of confidence across different inputs Si et al. (2023); Xiong et al. (2024). This results in an input bias, where some outputs sampled using the policy model will be assigned higher weights than others. Figure 2 shows the weight distribution of sampled outputs based on prompts from Ultrafeedback and the Mistral-sft-beta model, in which we observe significant variability in $w(x,y)$. To address this and ensure equal weighting of these outputs, we propose to align the weights in WPO.
@@
Greedy alignment. Here, we adjust the weights with respect to greedy decoding. Specifically, we adjust weights based on the maximum token probability among the set of all tokens in the subsequence, defined as:

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a promising solution to align large language models (LLMs) more closely with human values. Off-policy preference optimization, where the preference data is obtained from other models, is widely adopted due to its cost efficiency and scalability. However, off-policy preference optimization often suffers from a distributional gap between the policy used for data collection and the target policy, leading to suboptimal optimization. In this paper, we propose a novel strategy to mitigate this problem by simulating on-policy learning with off-policy preference data. Our Weighted Preference Optimization (WPO) method adapts off-policy data to resemble on-policy data more closely by reweighting preference pairs according to their probability under the current policy. This method not only addresses the distributional gap problem but also enhances the optimization process without incurring additional costs. We validate our method on instruction following benchmarks including Alpaca Eval 2 and MT-bench. WPO not only outperforms Direct Preference Optimization (DPO) by up to 5.6% on Alpaca Eval 2 but also establishes a remarkable length-controlled winning rate against GPT-4-turbo of 76.7% based on Gemma-2-9b-it. We release the code and models at https://github.com/wzhouad/WPO.
@@ 1 Introduction
We evaluate WPO on instruction following benchmarks, including Alpaca Eval 2 Li et al. (2023) and MT-bench Zheng et al. (2023). In the off-policy setting based on Ultrafeedback Cui et al. (2023), WPO improves the length-controlled winning rate against GPT-4-turbo on Alpaca Eval 2 by up to 14.9% over SFT model, outperforming DPO by up to 5.6%. Particularly, in the hybrid RL setting where the off-policy preference data is further enriched with on-policy outputs, WPO (Figure 1) achieves a new SOTA length-controlled winning rate of 76.7% on Alpaca Eval 2. Additionally, we find that WPO can be integrated into other loss functions for preference optimization and shows consistent improvements. Furthermore, we systematically compare the model performance in different RL settings. Our analysis reveals that the hybrid setting, which utilizes both on-policy and off-policy preference data, achieves the best results, and on-policy, dispreferred data is more important for preference optimization.
@@ 3.3 Weight Alignment
The objective of our weighting strategy is to simulate on-policy RL, where outputs are weighted according to how closely they align with on-policy behavior. For outputs generated by the current policy model, we expect their weights to be uniformly 1, while outputs that deviate from this on-policy behavior should receive smaller weights. However, due to the varying levels of confidence that LLMs exhibit across different inputs Si et al. (2023); Xiong et al. (2024), even outputs generated by the policy model may sometimes be assigned low weights. This introduces an unintended bias where some on-policy outputs receive lower weights purely because of lower model confidence based on the input, disrupting the uniformity we aim to achieve. Figure 2 shows the weight distribution of sampled outputs based on prompts from Ultrafeedback and the Mistral-sft-beta model, in which we observe significant variability in $w(x,y)$. To address this and ensure equal weighting of these outputs, we propose to align the weights in WPO.
@@
Greedy alignment. In this approach, we adjust the weights based on greedy decoding by comparing the probability of the current token with that of the most probable token in the vocabulary. Specifically, we adjust weights based on the maximum token probability among the set of all tokens in the subsequence, defined as:
@@ 4.2 Main Results and Ablation
Better base and reward models yield stronger results. To further enhance our model, we investigate using better base models and reward models. Specifically, we adopt the Gemma-2-9b-it Team et al. (2024) as the base model. In a setup similar to our hybrid approach, we sample five outputs from Gemma and one additional output from gpt-4-turbo. To rank these outputs, we apply ArmoRM Wang et al. (2024a, b) and use the best and worst outputs to form preference pairs. We use the same set of training hyperparameters used in Llama-3-Instruct. The model finetuned using WPO achieves a length-controlled win rate of 76.7% and a win rate of 77.8% on Alpaca Eval 2, demonstrating the effectiveness of this approach.